The solubility of Mg(OH)2 in pure water is 9.57 x 10^-3 g L^-1 . Its solubility (in g L^-1 ) in 0.02 moles Mg(NO3)2 will be

  1. The solubility of Mg(OH)2  in pure water is 9.57 x 10^-3 g L^-1  . Its solubility (in g L^-1 ) in 0.02 moles Mg(NO3)2 will be

Grade:9

1 Answers

nigamananda mishra
6 years ago
Solubility of Mg(OH)2 in pure water = 9.57 x 10-3 g/L
 = 9.57 x 10-3   /molar mass  mole/L
=  9.57 x 10-3   /58  mole/L = 1.65 x 10-4 mol/L
 
Kfor Mg(OH)2 = s x (2s)2 = 4s3 
= 4 x (1.65 x 10-4 mol/L)
= 17.96 x 10-12
 
Suppose, the solubility of Mg(OH)in the presence of  Mg(NO3)2  = s'
 
So, Mg2+ = s'+c = s' + 0.02
[OH'] = 2s'
 
Therefore, K= (s' + 0.02) (2s')2
 
17.96 x 10-12 =  4 (s' + 0.02) (s')2
 
or, 17.96 x 10-12/4 = s'3 + 0.02 (s' )2 , on neglecting s'3
 
4.4921 x 10-12 = 0.02 (s' )2
 
(s' )= 14.98 x 10-6 mol/l
 
Solubility in g/l
 = 14.98 x 10-6 x58 = 8.69 x 10-4 g/l